# Can AI Be as Creative as Humans?

### 03 Jan 2024 v1

@@ 4.3 Measuring Statistical Creativity of Large Language Model
The corollaries presented above provide an expansion of our prior theorems, illustrating the ways in which our creativity framework can be extended and reimagined with the introduction of a prompting setup. This expansion not only enhances our framework but also lays the groundwork for a deeper understanding of the behavior of large language models.
@@ 5 Achieving Statistical Creativity through Training
The discussions in the preceding sections have focused on establishing framework to determine whether trained models exhibit statistical creativity. This leads to a pivotal question: How can we steer a model towards achieving creativity? This section aims to crystallize and elucidate such an approach.
@@
Corollary 3 establishes that $L_{\bar{z}}[q_{S}]$ is constrained by $2\delta^{-1}E[q_{S}]+\mathcal{O}(\sqrt{\frac{Q(\delta/2)}{n}})$ (alternatively, $\frac{2}{\delta r_{\min}}\tilde{E}[q_{S}]+\mathcal{O}(\sqrt{\frac{\delta^{-2}Q(\delta/2)}{n}})$), where the first term minimizes with the training objective $E[q_{S}]$ (or $\tilde{E}[q_{S}]$), and the latter by increasing the sample size $n$. This corollary underscores that to achieve statistical creativity, one can minimize training objectives like $E[q_{S}]$ or $\tilde{E}[q_{S}]$ while simultaneously augmenting the training sample size $n$.
@@ 7 Conclusion
In this paper, we have embarked on an in-depth exploration of creativity within the realm of artificial intelligence. Our journey began with the introduction of Relative Creativity, a paradigm-shifting concept that redefines creativity assessment in AI systems from a comparative standpoint, moving away from absolute standards. This approach not only recognizes the inherent subjectivity in creative processes but also cleverly integrates it, drawing inspiration from the Turing Test’s comparative method of assessing intelligence. Progressing further, we introduced the concept of Statistical Creativity, which is instrumental in bridging the gap between theoretical constructs and empirical evaluation. By focusing on whether AI can mirror the creative output of specific human groups, this concept enables the quantifiable assessment of AI creativity and enhances the practical applicability of our theoretical framework. A significant stride was made with the application of these concepts to autoregressive models. We developed a practical measure for assessing statistical creativity in these models, particularly with next-token prediction. This measure’s adaptability to the prompting paradigm ensures its relevance and applicability to contemporary AI models, illustrating our commitment to keeping pace with technological advancements. Moreover, the introduction of Statistical Creativity Loss marks a cornerstone in our theoretical development. This concept provides invaluable insights for optimizing training objectives and dataset sizes, serving as a guidepost for achieving statistical creativity within a robust theoretical framework. In conclusion, our contributions lay a solid foundation for future explorations into AI creativity. We have not only advanced the theoretical discourse but also provided practical tools and methodologies for assessing and enhancing AI’s creative potential. Our work stands as a testament to the dynamic nature of AI research, and we aspire for it to act as a guiding beacon in the ongoing discourse on AI creativity, encouraging relative evaluations and empirical research. This paper, therefore, serves as both a theoretical contribution and a practical guide for those seeking to harness and understand the creative capacities of AI models.
@@
Limitations. This paper primarily focuses on conceptualizing Relative Creativity and proposing a methodology to assess the extent to which an AI model achieves this type of creativity. The development of a comprehensive dataset and the establishment of a benchmark for creativity evaluation are forthcoming endeavors. Additionally, we will test current advanced AI models to evaluate their creative performance. As we progress, we anticipate integrating our methods into the existing evaluation benchmark toolkit, align with the evolving landscape of AI technologies and assessment.

### 12 Jan 2024 v3

@@ 4.3 Statistical Creativity for Large Language Model
The corollaries presented above provide an expansion of our prior theorems, illustrating the ways in which our creativity framework can be extended and reimagined with the introduction of a prompting setup. This expansion not only enhances our framework but also lays the groundwork for a deeper understanding of the creativity of large language models.
@@ 5 Achieving Statistical Creativity through Training
The discussions in the preceding sections have focused on establishing framework to determine whether AI models exhibit statistical creativity. This leads to a pivotal question: How can we steer a model towards achieving creativity? This section aims to crystallize and elucidate such an approach.
@@
Corollary 3, combined with the Statistical Creativity Theorem (Theorem 1), reframes the discussion of AI’s creativity as a matter of its capacity to properly fit extensive data of existing human creators. If a model effectively fits the sufficient amount of data by minimizing creativity loss, it will generalize well on unseen data, according to Equation (2). Specifically, such a model is capable of generating creative works by using only information about its creators, thus exhibiting a level of creativity comparable to that of the creators themselves.
@@ 7 Conclusion
In this paper, we have embarked on an exploration of creativity within the realm of artificial intelligence, aiming to address the question of whether AI can be considered creative. We prove in theory that AI can indeed be considered as creative as humans under the mild condition that if AI can successfully personify existing human creators through fitting and learning their creations data. To reach this conclusion, we began our journey with the introduction of Relative Creativity, a paradigm-shifting concept that redefines creativity assessment in AI systems from a comparative standpoint, moving away from absolute standards. This approach not only recognizes the inherent subjectivity in creative processes but also cleverly integrates it, drawing inspiration from the Turing Test’s comparative method of assessing intelligence. Progressing further, we introduced the Statistical Creativity, focusing on whether AI can mirror the creative output of specific human groups. This concept facilitate the theoretical analysis of AI’s creativity. Besides, it enables the quantifiable assessment of AI creativity, enhancing the empirical applicability of our theoretical framework. Furthermore, in alignment with recent advancements in large language models (LLMs), we made progress by applying these concepts to autoregressive models. We developed a practical method for assessing statistical creativity in these models, particularly with next-token prediction. This measure’s adaptability to the prompting paradigm ensures its applicability to contemporary AI models. Moreover, the Statistical Creativity Loss provides insights for optimizing training objectives and dataset sizes, serving as a guidepost for achieving statistical creativity within a robust theoretical framework. In conclusion, our contributions lay a solid foundation for future explorations into AI creativity. We not only theoretically answer whether AI can be as creative as humans but also provided practical tools and methodologies for assessing and enhancing AI’s creative potential. Our work stands as a testament to the dynamic nature of AI research, and we aspire for it to act as a guiding beacon in the ongoing discourse on AI creativity, encouraging relative evaluations and empirical research. This paper, therefore, serves as both a theoretical contribution and a practical guide for those seeking to harness and understand the creative capacities of AI models.
@@
Limitations. This paper reduces the question of the AI’s relative creativity to the question of its ability to fit the massive amount of data. However, it does not provide an answer to the latter question. The possibility remains that AI’s inability to fit the required amount of data could impede its attainment of relative creativity. We would like to leave the discussion of it in future work. Besides, this paper conceptualize relative creativity and proposing a methodology to assess the extent to which an AI model achieves this type of creativity. The development of a comprehensive dataset and the establishment of a benchmark for creativity evaluation are forthcoming endeavors. Additionally, we will test current advanced AI models to evaluate their creative performance. As we progress, we anticipate integrating our methods into the existing evaluation benchmark toolkit, align with the evolving landscape of AI technologies and assessment..